A perfectly reflecting flat mirror is uniformly il
Answer #1

In case of perfect relection

Force exerted by beam is given as

F=2A\frac{I}{c}

thus

I=\frac{Fc}{2A}= \frac{1.29*10^{-6}*3*10^8}{2*2.61}=74.138W/m^2